CHAPTL13 ASSIGNMENT Question 43 of 50 > -/1 Prepare for under the following demandent mort red. Do not indien View 43 The shaCHAPTER 13 ASSIGNMENT < Question 43 of 50 > 0 Qu Todo Media Du LA
Answer #1

Hi
Hope it's 5000 shares, if not let me know. I will the answer accordingly.

Debit Credit
a Cash 65000
      Common Stock 25000 =5000*5
      Paid-in Capital in Excess of par value-Common Stock 40000
b Cash 65000
      Common Stock 25000 =5000*5
      Paid-in Capital in Excess of stated value-Common Stock 40000
c Organization Expense 65000
      Common Stock 25000 =5000*5
      Paid-in Capital in Excess of par value-Common Stock 40000
d Land 65000
      Common Stock 25000 =5000*5
      Paid-in Capital in Excess of par value-Common Stock 40000